### v1

# Synthesizing Programmatic Reinforcement Learning Policies with Large Language Model Guided Search

Abstract
Programmatic reinforcement learning (PRL) has been explored for representing policies through programs as a means to achieve interpretability and generalization. Despite promising outcomes, current state-of-the-art PRL methods are hindered by sample inefficiency, necessitating tens of millions of program-environment interactions. To tackle this challenge, we introduce a novel LLM-guided search framework (LLM-GS). Our key insight is to leverage the programming expertise and common sense reasoning of LLMs to enhance the efficiency of assumption-free, random-guessing search methods. We address the challenge of LLMs’ inability to generate precise and grammatically correct programs in domain-specific languages (DSLs) by proposing a Pythonic-DSL strategy – an LLM is instructed to initially generate Python codes and then convert them into DSL programs. To further optimize the LLM-generated programs, we develop a search algorithm named Scheduled Hill Climbing, designed to efficiently explore the programmatic search space to consistently improve the programs. Experimental results in the Karel domain demonstrate the superior effectiveness and efficiency of our LLM-GS framework. Extensive ablation studies further verify the critical role of our Pythonic-DSL strategy and Scheduled Hill Climbing algorithm.
1 Introduction
Deep reinforcement learning (DRL) has achieved great success from beating the world champion in Go [51] to powering the frontier natural language assistants [42, 3], and demonstrated great potential in robotics [28], autonomous vehicles [62], and recommendation systems [11]. However, approximating a policy using a deep neural network makes the decision-making process a black box and, therefore, less interpretable and trustable to human users [25]. Subsequently, carefully designed tools and costly human intervention are often required to deploy trustworthy DRL systems [18]. Moreover, DRL frequently encounters substantial performance declines when applied to previously unseen scenarios [31, 16], revealing another aspect of its challenges in generalization.
Recent programmatic reinforcement learning (PRL) methods have explored representing an RL policy using a program [2, 58, 5, 52]. Instead of learning a state-to-action mapping as in DRL, PRL synthesizes programs written in Domain-Specific Languages (DSLs) as human-readable policies that can be parsed and executed, making human inspection possible as an extra safety guard. Moreover, such structured program policies are shown to be able to capture high-level task-solving ideas, allowing for generalizing to a wide range of task variants [57]. Despite the encouraging results, the state-of-the-art PRL algorithms are notoriously inefficient and require tens of millions of program execution in environments [57, 37, 10]. Under the hood, they are search or RL algorithms without any assumption to the targeting problems. On the one hand, this allows generalization to all kinds of problems; on the other hand, the search time grows exponentially with increasing DSL complexity, making it intolerable for any practical use case.
Our key insight is that there is likely only a limited set of problems corresponding to specific program policies that are of human interest – hence, we can utilize reasonable assumptions to prune the space. Recently, Large Language Models (LLMs) have been demonstrated to possess internet-scale knowledge that can be retrieved by a natural language interface [61, 65, 56]. If we view the text on the internet as a “text projection” of human civilization, an LLM as a knowledge base should be able to provide hints to “prune” the program search paths that are out of the “human interest scope.” With this intuition, we conjecture LLMs can be utilized to bootstrap the sample efficiency of search-based PRL algorithms, pushing PRL one step closer to practical adoption.
To this end, we aim to develop a PRL framework that utilizes LLMs to produce task-solving program policies while minimizing the number of program executions in environments. Directly instructing LLMs to synthesize DSL programs for solving PRL tasks faces three fundamental challenges: (1) LLMs may lack the domain knowledge of the PRL tasks, e.g. environment dynamics, what an agent can obverse or how it can act, (2) the training data of modern LLMs are mostly natural language texts and general-purpose programming languages, e.g. Python and C++, which can be quite different from the DSLs used in PRL, and (3) there is no apparent mechanism for optimizing LLMs to produce programs that maximize rewards since the best-performing LLMs are privately owned.
To combat these challenges, we present an LLM-guided search (LLM-GS) framework leveraging the programming skills and common sense of LLMs and the effectiveness of search algorithms. (1) To familiarize LLMs with PRL tasks, we devise domain and task-aware prompts that convey PRL domain knowledge to LLMs while avoiding leaking task-solving information. (2) To mitigate the gap between general-purpose programming languages and DSLs, we design a Pythonic-DSL strategy that allows LLMs to generate more precise and grammatically correct DSL programs by first producing Python programs. (3) To further optimize the LLM-generated programs, we propose a search algorithm, Scheduled Hill Climbing (Scheduled HC), to efficiently improve programs.
We compare our proposed LLM-GS framework in the Karel domain to various existing PRL methods [57, 37, 10]. The experimental results demonstrate that LLM-GS is significantly more effective and efficient than the baselines. Extensive ablation studies show that (1) our proposed Pythonic-DSL strategy leads to a higher ratio of executable (i.e. grammatically correct) programs and a higher average return compared to directly generating DSL programs, (2) our proposed search method Scheduled HC achieves the best efficiency among existing search algorithms, (3) initializing the search population of Scheduled HC using LLM-generated programs is significantly more efficient than randomly sampled programs.
2 Related work
Programmatic reinforcement learning (PRL). PRL represents RL policies using more structured and potentially more interpretable and generalizable representations, such as decision tree [5], state machine [27, 32, 36], symbolic expression [58, 59, 6, 33], Logic programming language [29], and program written in domain-specific language (DSL) [2, 52, 54, 66, 46, 39, 38, 57, 37, 10]. Trivedi et al. [57], Liu et al. [37] devised PRL tasks in the Karel domain [44] and proposed learning embedding spaces of programs using variational autoencoders [30] and then optimizing program embeddings using cross-entropy method or RL. Carvalho et al. [10] achieve the state-of-the-art results in the Karel domain [44] by searching in a programmatic space of AST structure. Despite the encouraging results and the generality of these methods, they are notoriously inefficient, requiring tens of millions of program executions to obtain task-solving programs. In this work, to devise an efficient PRL framework, we make the first attempt to leverage the knowledge and the reasoning ability of LLMs to generate a set of initial programs to bootstrap search algorithms; moreover, we propose Scheduled HC, a search algorithm that best utilizes the LLM-generated programs.
Large language models for code generation. With a remarkable ability to understand and generate natural languages and codes, LLMs have been widely adopted for code generation [7, 1, 40, 64, 48] via neural program synthesis methods [35, 53, 17, 47, 63, 34, 43, 4]. These works target general-purpose programming languages for software development, such as Python and C++, with abundant data on the internet; in contrast, we aim to solve PRL tasks that require writing programs restricted by given domain-specific languages. Prior works exploring using LLMs to synthesize DSL programs via providing LLMs with DSL grammars and few-shot examples, hindsight relabeling, and prioritized experience replay [60, 21, 9]. While these works focus on string and array transformation, and abstract reasoning [15], we leverage LLMs to synthesize DSL program policies to be executed in an RL environment and maximize return.
Search-based program synthesis. Various search algorithms have been developed for program-by-example (PBE) [22, 20, 45, 23, 43, 4], whose goal is to find programs that satisfy given examples, e.g. input/output string pairs. Recent works have explored utilizing search algorithms in DSL to learn libraries [19, 21]. In the regime of programmatic RL, Carvalho et al. [10] recently applied the Hill Climbing algorithm to the Karel benchmark, achieving state-of-the-art performance. Instead of randomly sampling programs via search algorithms, we present a framework that integrates knowledge from an LLM with a search algorithm to significantly improve the sample efficiency.
{mdframed}
[font=]
$\text{Program}\ \rho$ $\coloneqq\text{DEF}\ \text{run}\ \text{m}(\ s\ \text{m})$
$\text{Repetition}\ n$ $\coloneqq\text{0..19}$
[…]
$\text{Statement}\ s$ $\coloneqq\text{WHILE}\ \text{c}(\ b\ \text{c})\ \text{w}(\ s\ \text{w})\ |\ s_{1}s_{2}\ |\ a\ |$
[…]
$\ \text{IFELSE}\ \text{c}(\ b\ \text{c})\ \text{i}(\ s_{1}\ \text{i})\ \text{ELSE}\ \text{e}(\ s_{2}\ \text{e})$
Figure 1: The Karel DSL grammar. It describes the Karel domain-specific language’s actions, perceptions, and control flows. The domain-specific language is obtained from Liu et al. [37].
Figure 2: An example Karel task – DoorKey. The agent first needs to find the key (marker) in the left room, which will open the door (wall) to the right room. Navigating to the goal marker in the right room and placing the picked marker on it will grant the full reward for the task. This sparse-reward task has been found to pose significant challenges to previous PRL methods, as it necessitates a greater capability in long-horizon strategy formulation.
3 Preliminary
The Karel domain. We first review the Karel domain, the de facto test bed for programmatic reinforcement learning research [8, 12, 50, 24, 13]. The Karel domain-specific language illustrated in Figure 2 is a robot programming language to control the Karel agent in a 2D grid world. The agent’s actions include moving as well as interacting with the environment by picking up and putting down objects (markers). The perceptions check for obstacles and markers, which allows observing the environment. Lastly, control flows, e.g. REPEAT, WHILE, IF, and IFELSE, enable describing complex decision-making logics.
We illustrate a Karel task, DoorKey, in Figure 2. The agent needs to explore the left room using the DSL actions such as move and turnLeft as well as the perceptions such as frontIsClear and markersPresent to find the key (marker). After the key is picked up via pickMarker, the wall to the right room will unblock. The agent then needs to place another marker on the goal marker, which appears in the right room, to receive the full reward for the task. Appendix C presents all Karel tasks.
The state-of-the-art performance in Karel is achieved by iteratively searching for improved programs according to their episodic return using search algorithms [10]. Next, we introduce the concept of search space, which defines how search algorithms find neighbor programs from a set of current candidates for improved performance, and then dive into the search algorithm details.
Programmatic space vs. latent space. In the programmatic space, a program $P$ can be represented as an abstract syntax tree (AST), where each leaf node in the AST represents a program token [10]. To obtain its neighborhood program, a node is sampled from its AST, and then replaced by a subtree randomly generated using the DSL’s production rules and sampling strategies. Another choice is to search in a learned latent space [57, 37], constructed by training a variational autoencoder (VAE) on randomly generated DSL programs. To sample a neighbor program, a small noise is added to the latent embedding of the current program candidate and the VAE decoder can generate a program from the noise-corrupted embedding. Appendix B contains more details of search spaces.
Hill climbing (HC). This search algorithm climbs whenever there is a higher place, i.e., moves to a program having a higher episodic return. Given a program, HC generates its $k$ neighborhood programs from the search space. These programs are evaluated in the environment. If one of them has a higher episodic return than the initial program, it is set as the search center program, and then this process repeats. Otherwise, the algorithm halts and returns the best program evaluated.
Cross-entropy method (CEM). CEM [49] first generates $k$ neighborhood programs from its latent space and evaluates their episodic return on the Karel environment [57]. Next, if the average return over the top-$E$ neighborhood programs is higher than the best-seen average return, it continues the process on the mean latent embedding of the top-$E$ neighborhood programs. When the average return no longer increases, the program with the highest episodic return is returned.
Cross-entropy beam search (CEBS). CEBS extends CEM to maintain a set of $E$ candidate programs to perform beam search. In other words, CEBS searches all the neighborhoods of the top-$E$ programs.
Note that HC can search in both programmatic and latent spaces, while CEM and CEBS are designed for continuous search spaces, i.e., latent program space only. All of these search algorithms share the same weakness – they search surrounding programs in a narrow proximity thus the distance between the initial program and the optimal program directly upper-bounds the algorithm’s sample efficiency.
Figure 3: Large language model-guided search (LLM-GS). (a) With task description and the Pythonic-DSL instruction, LLM generates Python programs that are subsequently converted to DSL programs. (b) These initial programs serve as the initial population of our proposed Scheduled Hill Climbing, which evaluates the episodic return of the neighboring programs to update the current candidate program with increasing neighborhood size over search steps.
4 Large language model-guided search for efficient programmatic reinforcement learning
Table 1: Activate LLM’s task awareness. To bridge the domain gap between the target task and the LLM’s knowledge, we craft a carefully curated prompt with relevant knowledge of solving the task without explicitly dictating the specific programming approach or implementation. Take the task DoorKey here as an example: given task name, map description, initial position, and task goal, the LLM is encouraged to solve the specified task with its programming skills and common sense. More details can be found in Section E.2.
Knowledge Prompt Text
Task Name DoorKey
Map Description The map is a $8\times 8$ grid surrounded by walls that is vertically split into two chambers. The left chamber is $6\times 3$ grid and the right chamber is $6\times 2$ grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Initial Position The agent starts on a random cell on the left chamber facing east.
Task Goal The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
The hill climbing algorithm (HC) is the state-of-the-art programmatic reinforcement learning (PRL) approach in Karel, at the cost of tens of millions of program interactions [10]. This prevents its application to real-world decision-making problems, where program-environment interaction at this scale is inapplicable. Hence, we aim to maintain the high episodic return while reducing programs executed in the environments, i.e., improves the sample efficiency of PRL. To this end, our key insight is to utilize a large language model (LLM), hypothesizing that its abundant world knowledge, including programming skills and common sense, may bootstrap the inherently assumption-less, random-guessing HC. However, to implement this idea, many challenges rooted in the environment’s domain gap (Section 4.1), DSL’s language barrier (Section 4.2), and the inability to optimize closed-source LLMs (Section 4.3) must be addressed.
4.1 Domain and task-aware prompting
An inexperienced LLM user might directly ask an LLM: “Write a program to solve the Karel task DoorKey,” since this is how average users work with LLM-based software development assistants. Obviously, repeating this until the LLM spits out a correct program will not be more efficient than existing search algorithms. The first challenge we encountered was that the LLM lacks background knowledge of the PRL task at hand. After all, there is no guarantee that the LLM has seen or learned the text or code content related to the specific PRL task. Therefore, it is crucial to describe the task and the environment to the LLM in detail using natural language so that it makes the right assumptions based on its learned common sense. Here, we devise a prompting strategy that converts the PRL task’s domain knowledge into natural language sentences that LLMs can process and reason. Table 1 shows the prompt we carefully wrote for Karel DoorKey. Note how the instructions humanly describe the task but do not leak the actual programming skills to implement the required navigation. Now that the LLM possesses minimal knowledge of the task, we may further instruct it to generate programs that solve the desired task.
4.2 Generating DSL programs with Pythonic-DSL strategy
Given LLMs’ widespread success in assisting software development, one might assume an LLM can naively generate DSL programs. Yet, precisely generating DSL programs turns out to be quite challenging for LLMs – their training data typically consists of natural language corpus and general-purpose programming language codes, thus specific DSLs used in PRL tasks are actually quite exotic for them. Note that this language barrier is beyond lexical syntax differences. For example, the DSL may specify rules that limit the usage of temporary variables constrained by the actual robotic hardware, which is rarely a concern for common Python / JavaScript code found on the internet.
Despite the issue, we still hope to leverage the best of LLM’s programming skills to synthesize DSL programs. To this end, we derive a solution from the following assumptions: (1) The LLM is proficient in a general-purpose programming language, such as Python, (2) The DSL can be represented by a possibly restricted version of the general-purpose programming language, and (3) The LLM understands the restriction in natural language, if any. Our Pythonic-DSL strategy instructs the LLM to generate Python programs instead, given the Karel rules and constraints we wrote in English, and then later convert it into the Karel DSL. As shown in Table 2, DSL details, including the action, perception functions, and language constraints, are provided. [^1] Additionally, task-agnostic environment knowledge is also given. See Appendix F for the full prompt. Thus, the LLM can generate programs based on its innate general programming skills and its understanding of the DSL description. One caveat is that the LLM still occasionally generates Python programs that cause Python-to-DSL parsing failures. An empirical useful mitigation is instructing the LLM to generate the DSL program converted from its own output Python program as a backup. At this point, we have developed the mechanism to instruct an LLM to generate DSL programs given the PRL task and the DSL grammar in natural language.
Table 2: The Pythonic-DSL instruction. The Karel DSL is specified via a constrained version of Python with pre-defined functions including Pythonic-DSL perceptions that allow the agent to observe, and Pythonic-DSL actions that enable the agent to interact. This Pythonic-DSL description is fed into the LLM via the system prompt so that it can generate Python code that can later be converted into Karel DSL. We will show that this approach outperforms the direct generation of DSL or Python programs in Table 3. The full prompt is presented in Section E.1.
Knowledge Prompt Text
Pythonic-DSL frontIsClear(): Returns True if there is no wall in front of the agent.
Perceptions markersPresent(): Returns True if there exist markers on the current cell.
(more perceptions omitted)
Pythonic-DSL move(): Asks the agent to move forward one cell … (truncated)
Actions turnLeft(): Asks the agent to rotate 90 degrees counter-clockwise.
pickMarker(): Asks the agent to pick up one marker from the current cell.
(more actions omitted)
Language Constraints - do not define other functions besides run()
- do not define variables
(more rules omitted)
4.3 Optimizing the program policy with Scheduled Hill Climbing
Although the aforementioned techniques can already instruct the LLM to produce DSL programs that solve the target task to a certain degree, the episodic return can still be far from optimal. Since best-in-class LLMs are all proprietary APIs with no fine-tuning service; therefore, conventional gradient-based policy optimization is out of the question. To further optimize the program generated by the LLM, we explore initializing search populations of search algorithms using LLM-generated programs. Via extensive experiments with the HC, CEM, and CEBS search algorithms on both program space and latent space whenever applicable, we discover a key improvement that is crucial when initializing search from LLM-generated programs – program-environment interaction scheduling for a more efficient allocation of the interaction budget. The intuition is that LLMs can often provide good initialization and therefore it makes sense to keep a small search budget, i.e. with a smaller population. After a while, if an optimal program is not yet found, we can gradually increase the budget and facilitate broader exploration. We design a scheduler based on this intuition:
$\log_{2}{k(n)}$ $=(1-r(n))\log_{2}{K_{start}}+r(n)\log_{2}{K_{end}}\text{,}$ (1)
where $n$ represents the number of evaluated programs, $k(n)$ denotes a function indicating the current number of neighborhood programs based on the number of evaluated programs $n$. This function’s logarithm is a linear combination of the logarithms of two hyperparameters, $K_{start}$ and $K_{end}$, which signify the initial and terminal numbers of neighbors to search, respectively. The variable $r(n)$ governs the linear ratio, which itself is a sinusoidal function that gradually increases from 00 to $1$ throughout the evaluation of a total of $N$ programs (see Appendix A). With this scheduler, the number of neighbors $k$ is a function of the execution budget used, growing from $K_{start}$ to $K_{end}$. We apply this scheduler to Hill Climbing and call this method Scheduled Hill Climbing (Scheduled HC).
Overall, our best recipe consists of the following steps. First, we sample DSL programs for the LLM using the aforementioned techniques. Next, these programs are evaluated in the environment for the episodic return. In some easy tasks, some generated programs might have achieved perfect returns and thus are optimal programs. If no optimal program is found, we sort the programs based on the total reward decreasingly. Next, programs are selected from the sorted list, each serves as an initial program for the HC search, which is the best-performing search algorithm from our extensive experiments. HC searches $k$ neighbors at each step following the scheduler in Eq. 1. This process is repeated until either the optimal program is found or we meet the maximum program-environment interaction budget allowed ($N$). Figure 3 presents an overview of our framework.
(a) StairClimber
(b) Maze
(c) FourCorners
(d) TopOff
(e) Harvester
(f) CleanHouse
(g) DoorKey
(h) OneStroke
(i) Seeder
(j) Snake
Figure 4: Efficiency in the Karel domain. We compare our proposed LLM-guided search (LLM-GS) framework against existing methods, LEAPS, HPRL, CEBS, and HC in the Karel and Karel-Hard problem sets. The results show that our LLM-GS is significantly more efficient than these methods.
5 Experiments
5.1 Evaluation setup
PRL tasks. We evaluate our proposed framework LLM-guided search (LLM-GS) using the Karel tasks from the two problem sets: Karel [57] (StairClimber, Maze, FourCorners, TopOff, Harvester, and CleanHouse) and Karel-Hard [37] (DoorKey, OneStroke, Seeder, and Snake). More details of the task can be found in Appendix C.
Baselines. We compare our proposed framework with the following methods: LEAPS [57], HPRL [37], CEBS [10], and Hill Climbing [10] (the current state-of-the-art). More details of these methods are presented in Appendix D.
Metrics. For a fair comparison of the sample efficiency, we use the number of programs evaluated to represent the program-environment interaction budget. A method is a more sample-efficient PRL algorithm if it achieves a higher average episodic return on a fixed budget. Specifically, for each task, the number of task variances is $C$. One program evaluated means a program is executed on all $C$ task variances to obtain the average return. We set the number of task variances $C=32$, the maximum number of program evaluation $N=10^{6}$, i.e., the interaction budget.
Setup. We use gpt-4-turbo [1] (temperature=1.0, top_p=1.0) as our LLM module to generate the initial search population with 32 programs. The scheduler of our proposed Scheduled HC starts from $K_{start}=32$ to $K_{end}=2048$. We evaluate our LLM-GS and HC with $32$ random seeds and $5$ seeds for LEAPS, HPRL, and CEBS.
5.2 LLM-guided search significantly improves the sample efficiency
Figure 4 presents our full experimental result, indicating that our proposed LLM-GS surpasses all the existing methods by a large margin on almost all the tasks. On the Karel set, our framework completely solves these tasks and exhibits the best sample efficiency. Also, LLM-GS performs extraordinarily well on Seeder, OneStroke, and DoorKey, which are among the hardest of the Karel-Hard set. We highlight our DoorKey result, with an efficiency improvement from not even converging at 1M to converge within around 50K. Existing PRL algorithms struggle at DoorKey due to the fact that it is a two-stage task with a sparse reward function. Therefore, most of the baselines cannot escape the local maxima and converge around $0.5$. Curiously, our method starts at an episodic return of $\approx 0.5$ from the beginning and converges quickly without being trapped by the local maxima. To analyze, we examine one output program from LLM shown in Figure 5. The LLM clearly understands that DoorKey is a two-stage task and thus generates a corresponding two-stage structured program, which is much easier for later search to find the optimal program. The full dialogue is presented in Appendix E.
Figure 5: Example on DoorKey. This example shows how our search method improves an LLM-initialized program to an optimal one. The original program (left) has a two-stage structure but lacks navigation ability. The improved program (right) solves this by enhancing its navigating ability on both stages, allowing for solving the task.
5.3 Ablation studies
Which language to generate, Python or DSL? Do both!
To justify the effectiveness of our proposed Pythonic-DSL strategy, we conduct experiments on generating either Python or DSL programs with the same 8 seeds and present the results in Table 3. The acceptance rate calculates the ratio of executable DSL programs, while the best return refers to the highest episodic return among the legal programs. Our method achieves the highest acceptance rate on 9 / 10 tasks and the highest best return on 7 / 10 tasks. The complete prompts are presented in Appendix F.
Table 3: Pythonic-DSL strategy ablation. We compare the programs generated using our proposed Pythonic-DSL strategy to directly generated Python or DSL programs. Our method achieves the highest acceptance rate and best return, justifying the effectiveness of our Pythonic-DSL strategy.
Task Python DSL Pythonic-DSL (ours)
Acceptance Rate Best Return Acceptance Rate Best Return Acceptance Rate Best Return
StairClimber 88.02$\pm$3.571.00$\pm$0.00 45.31$\pm$11.590.75$\pm$0.65 88.54$\pm$4.541.00$\pm$0.00
Maze 72.40$\pm$7.490.94$\pm$0.08 57.81$\pm$8.120.88$\pm$0.07 92.45$\pm$3.890.98$\pm$0.05
FourCorners 83.07$\pm$5.041.00$\pm$0.00 93.49$\pm$2.431.00$\pm$0.00 94.01$\pm$2.431.00$\pm$0.00
TopOff 90.89$\pm$2.070.97$\pm$0.07 94.53$\pm$4.291.00$\pm$0.00 96.35$\pm$3.081.00$\pm$0.00
Harvester 67.45$\pm$7.510.98$\pm$0.03 79.69$\pm$7.640.98$\pm$0.03 80.21$\pm$5.510.95$\pm$0.09
CleanHouse 62.50$\pm$3.450.27$\pm$0.28 76.30$\pm$6.160.36$\pm$0.14 84.64$\pm$4.290.29$\pm$0.27
DoorKey 90.10$\pm$3.860.63$\pm$0.13 75.78$\pm$15.090.55$\pm$0.05 94.53$\pm$3.600.63$\pm$0.14
OneStroke 85.42$\pm$4.770.77$\pm$0.06 85.68$\pm$7.540.74$\pm$0.09 95.31$\pm$3.080.83$\pm$0.08
Seeder 73.70$\pm$4.770.93$\pm$0.14 88.28$\pm$4.880.66$\pm$0.10 74.74$\pm$7.100.90$\pm$0.14
Snake 87.50$\pm$2.760.09$\pm$0.00 30.47$\pm$9.020.09$\pm$0.00 98.96$\pm$1.040.09$\pm$0.00
Our proposed Scheduled HC is the most efficient search method. We compare various search methods in both programmatic and latent spaces in DoorKey and CleanHouse. In the learned latent space, we run CEM, CEBS, HC-$250$ (HC with a fixed population size of $250$), and our proposed Scheduled HC; in the programmatic space, we run HC-$k$$\in\{32,256,2048\}$ and our proposed Scheduled HC. The results presented in Figure 7 show that HC in the programmatic space achieves the best performance compared to latent space search. However, HC-$k$ with a fixed population size only specializes in specific tasks. To be specific, HC-$32$ for CleanHouse and HC-$2048$ for DoorKey. In contrast, our proposed Scheduled HC achieves comparatively the best performance in both tasks.
LLM provides a better initial search population. To verify the efficacy of using LLM-generated programs as search initialization for Scheduled HC, we compare it to using randomly sampled programs as initialization in DoorKey and CleanHouse. The results shown in Figure 7 suggest that initializing the search with LLM-generated programs significantly improves the sample efficiency.
(a) DoorKey
(b) CleanHouse
Figure 6: Comparing search algorithms. We compare the proposed Scheduled HC with latent space CEM, CEBS, HC-$250$, and Scheduled HC, and programmatic space HC-$k$$\in\{32,256,2048\}$. Our method achieves comparatively good efficiency in DoorKey and performs best in CleanHouse, justifying the efficacy of our proposed search method.
(c) DoorKey
(d) CleanHouse
Figure 7: Comparing initializations. We compare initializing the search population using LLM-generated programs or randomly sampled programs. The results show that starting from LLM-generated programs significantly improves efficiency and performance, especially in DoorKey, highlighting the effectiveness of our proposed framework.
5.4 Beyond initialization – a preliminary attempt on LLM revision
Given the encouraging empirical results from LLM-initialized programs, one might be curious whether LLM can keep revising and improving its programs once given feedback, e.g. reward, from the environment. For completeness, we additionally conduct studies with the following LLM-based revision:
  • Regeneration. Inspired by Chen et al. [14] and Olausson et al. [41], we instruct the LLM to re-generate non-duplicate programs with previously generated programs fed as inputs.
  • Regeneration with reward. In addition to the historical program list, the episodic return of each program is appended to the LLM’s inputs.
  • Agent execution trace. Inspired by prior arts that leverage environment knowledge [55, 61] and code repair [14, 41], we feed the program with the best episodic return and its execution traces into the LLM.
  • Agent and program execution trace. Inspired by Hu et al. [26], we additionally point out the line of the program currently being executed as extra hints to the LLM.
Figure 8: LLM revision. The performance gain of LLM revision quickly saturates within the first two rounds.
We conduct the LLM revision experiments on DoorKey for five revision rounds, each round with $32$ programs with $5$ seeds. The results are presented in Figure 8. “Agent and program execution trace” performs best, indicating the importance of feeding LLM as much information as possible. Yet, the improvement quickly saturates within the first two rounds, and no significant gain is observed with more revision rounds. This indicates that only using LLM to revise the program may not be sufficient to solve the task. On the other hand, our proposed Scheduled HC can converge to an optimal program – and is “free,” i.e. no expensive API calls or “API usage has reached your notification threshold” email.
6 Conclusion
We propose an LLM-guided framework (LLM-GS), that leverages the programming skills and common sense of LLM to bootstrap the efficiency of assumption-free, random-guessing search methods. We propose a Pythonic-DSL framework to address the inability of LLMs to generate precise and grammatically correct programs. To further maximize the programs generated by LLMs, we design a search algorithm, Scheduled Hill Climbing, which can explore the programmatic search space to consistently improve the programs. The evaluation result on Karel environment shows the effectiveness of our proposed LLM-guided framework. Extensive ablation studies verify the necessities of our proposed Pythonic-DSL and Scheduled Hill Climbing.
Appendix
[…]
Appendix A Schedule hill climbing detail
This Scheduler is designed for HC. $n$ is the number of programs evaluated, and $K_{start}$ and $K_{end}$ are hyper-parameters representing the initial and terminal number of neighbors to search, respectively. $r(n)$ is a sinusoidal function that smoothly increase from 00 to $1$ over the course of $N$ total programs evaluated in the environment. We provide detailed equation in Equation 2 and Equation 3
$\log_{2}{k(n)}$ $=(1-r(n))\log_{2}{K_{start}}+r(n)\log_{2}{K_{end}}\text{,}$ (2)
$2r(n)=\sin{\left[(\frac{2\log{n}}{\log{N}}-1)\times\frac{\pi}{2}\right]}+1\text{.}$ (3)
Appendix B Karel DSL details
There are five symbols in the Karel DSL: statement, condition, action, boolean, and number. The production rules and Python converting rules are listed in Table 4. The statement counts mean total usage of converting one statement into two statements at the root of the abstract syntax tree. And the statement counts are limited to six in search of optimal programs. The program length is the total number of tokens used in string representation. The token limit is set to be forty-four. The depth in DSL is the recursive call of control flows. The depth is limited to four. These limitations are the same as the original settings in LEAPS [57]. The probabilities of the production rules are listed in Table 6. The mutation process selects one random node in the abstract syntax tree (AST). We can use the production rules to sample a sub-tree to replace the original node. The Karel syntax has many symbol-related parentheses, e.g. “i(” and “i)” which are related to the “IF” and “IFELSE”, thus LLM sometimes makes mistakes translating Python to DSL. We list common mistakes LLM made in Table 5 and how we solve these problems.
Table 4: Python to DSL converting rules.
Python DSL
def run(): s DEF run m( s m)
while b: s WHILE c( b c) w( s w)
if b: s IF c( b c) i( s i)
if b: s else: s IFELSE c( b c) i( s i) ELSE e( s e)
for i in range(n): s REPEAT R=n r( s r)
not h not c( h c)
frontIsClear() frontIsClear
leftIsClear() leftIsClear
rightIsClear() rightIsClear
markersPresent() markersPresent
noMarkersPresent() noMarkersPresent
move() move
turnLeft() turnLeft
turnRight() turnRight
putMarker() putMarker
pickMarker() pickMarker
Table 5: These are the common mistakes LLM makes. We implement a simple program to correct these mistakes.
Type Before After
Brackets removal DEF run m( move() m) DEF run m( move m)
Brackets separation DEF run m(move m) DEF run m( move m)
Brackets addition DEF run m( WHILE frontIsClear w( move w) )m DEF run m( WHILE c( frontIsClear c) w( move w) )m
Brackets repairment DEF run m( move ) DEF run m( move m)
If to IFELSE DEF run m( IF c( frontIsClear c) i( move i) ELSE e( turnLeft e) )m DEF run m( IFELSE c( frontIsClear c) i( move i) ELSE e( turnLeft e) )m
Redundant symbols removal DEF run m( move() m m) DEF run m( move m)
Illegal symbols transformation DEF run m( WHILE c( True c) w( move w) ) DEF run m( REPEAT r=19 r( move r) m)
Table 6: Probabilities of the production rules.
Category Rule Probability
Program P Statement 1.0
Statement S While 0.15
Repeat 0.03
Concatenate 0.5
If 0.08
Ifelse 0.04
Action 0.2
Condition c Boolean 0.9
not 0.1
Action a move 0.5
turnLeft 0.15
turnRight 0.15
putMarker 0.1
pickMarker 0.1
Boolean b frontIsClear 0.5
leftIsClear 0.15
rightIsClear 0.15
markersPresent 0.1
noMarkersPresent 0.1
Number n i (0 $\leq$ i $\leq$ 19) 0.05
Appendix C Karel tasks
There are five knowledge we tell the Large Language Model: task name, task map, task agent position, task goal, and task return. Here is all the information about these ten tasks. Appendix F show how these five different components are filled into the placeholders in the user prompt. There are two sets of tasks: Karel [57] and Karel-Hard [37]. Section C.1 lists all six Karel tasks, and Section C.2 lists all four Karel-Hard tasks. There are six Karel tasks: StairClimber, FourCorners, TopOff, Maze, Harvester, and CleanHouse. Example figures of Karel tasks are in Figure 9. On the other hand, four of the Karel-Hard tasks are DoorKey, OneStroke, Seeder, and Snake. Figure 10(d) shows the initial and desired states of the Karel-Hard tasks.
C.1 Karel tasks
(a) StairClimber
(b) fourCorner
(c) TopOff
(d) Maze
(e) Harvester
(f) CleanHouse
Figure 9: Illustrations of the initial and desired final state of each task in the Karel Problem set introduced in by Trivedi et al. [57]. Note that these illustrations are from Trivedi et al. [57] except for StairClimber, TopOff, and FourCorners. We align our setting with Carvalho et al. [10] to evaluate these tasks in a map of 12x12. The position of markers, walls, and agent’s position are randomly set according to the configurations of each task.
C.1.1 The task prompt of StairClimber
Knowledge Prompt Text
Task Name StairClimberSparse
Map Description The map is a 12x12 grid surrounded by walls with stairs formed by walls and a marker is randomly initialized on the stairs as a goal.
Initial Position The agent starts on a random position on the stairs facing east.
Task Goal The goal of the agent is to reach a marker that is also randomly initialized on the stairs.
Task Reward If the agent reaches the marker, the agent receives 1 as an episodic return and 0 otherwise. If the agent moves to an invalid position, i.e. outside the contour of the stairs, the episode terminates with a -1 return.
C.1.2 The task prompt of Maze
Knowledge Prompt Text
Task Name MazeSparse
Map Description The map is a complex 8x8 grid surrounded by walls and a random marker is placed on an empty cell as a goal.
Initial Position The agent starts on a random empty cell of the map facing east.
Task Goal The goal of the agent is to reach the goal marker.
Task Reward If the agent reaches the marker, the agent receives 1 as an episodic return and 0 otherwise.
C.1.3 The task prompt of FourCorners
Knowledge Prompt Text
Task Name FourCorners
Map Description The map is an empty 12x12 grid surrounded by walls.
Initial Position The agent starts on a random cell on the bottom row of the map facing east.
Task Goal The goal of the agent is to place one marker in each corner of the map.
Task Reward Return is given by the number of corners with one marker divided by 4.
C.1.4 The task prompt of TopOff
Knowledge Prompt Text
Task Name TopOff
Map Description The map is a 12x12 grid surrounded by walls with markers randomly placed on the bottom row of the map.
Initial Position The agent starts on the bottom left cell of the map facing east.
Task Goal The goal of the agent is to place one extra marker on top of every marker on the map.
Task Reward Return is given by the number of markers that have been topped off divided by the total number of markers. Picking up the marker will terminate the episode with a -1 return.
C.1.5 The task prompt of Harvester
Knowledge Prompt Text
Task Name Harvester
Map Description The map is a 8x8 grid surrounded by walls that starts with a marker on each cell.
Initial Position The agent starts on a random cell on the bottom row of the map facing east.
Task Goal The goal of the agent is to pick up every marker on the map.
Task Reward Return is given by the number of picked-up markers divided by the total number of markers.
C.1.6 The task prompt of CleanHouse
Knowledge Prompt Text
Task Name CleanHouse
Map Description The map is a complex 14x22 grid made of many connected rooms and is surrounded by walls. There are ten markers randomly placed adjacent to the walls.
Initial Position The agent starts on a fixed cell facing south.
Task Goal The goal of the agent is to pick up every marker on the map.
Task Reward Return is given by the number of picked-up markers divided by the total number of markers.
C.2 Karel-Hard tasks
(a) DoorKey
(b) OneStroke
(c) Seeder
(d) Snake
Figure 10: Illustrations of the initial and final state of each task in the Karel-Hard problem set introduced in by Liu et al. [37]. The position of markers, walls, and agent’s position are randomly set according to the configurations of each tasks.
C.2.1 The task prompt of DoorKey
Knowledge Prompt Text
Task Name DoorKey
Map Description The map is a 8x8 grid surrounded by walls that is vertically split into two chambers. The left chamber is 6x3 grid and the right chamber is 6x2 grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Initial Position The agent starts on a random cell on the left chamber facing east.
Task Goal The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
Task Reward Picking up the first marker yields a 0.5 reward, and putting a marker on the goal marker yields an additional 0.5.
C.2.2 The task prompt of OneStroke
Knowledge Prompt Text
Task Name OneStroke
Map Description The map is given by an empty 8x8 grid surrounded by walls.
Initial Position The agent starts on a random cell of the map facing east.
Task Goal The goal of the agent is to visit every grid cell without repeating. Visited cells become a wall that terminates the episode upon touching.
Task Reward Return is given by the number of visited cells divided by the total number of empty cells in the initial state.
C.2.3 The task prompt of Seeder
Knowledge Prompt Text
Task Name Seeder
Map Description The map is given by an empty 8x8 grid surrounded by walls..
Initial Position The agent starts on a random cell of the map facing east.
Task Goal The goal of the agent is to place one marker in every empty cell of the map.
Task Reward Return is given by the number of cells with one marker divided by the total number of empty cells in the initial state.
C.2.4 The task prompt of Snake
Knowledge Prompt Text
Task Name Snake
Map Description The map is given by an empty 8x8 grid surrounded by walls with a marker randomly placed on the map.
Initial Position The agent starts on a random cell of the map facing east.
Task Goal The agent acts like the head of a snake, whose body grows each time a marker is reached. (No need to pick it up.) Every time a marker is reached, the body of the agent grows one marker. The goal of the agent is to touch the marker on the map without colliding with the snake’s body, which terminates the episode. Each time the marker is reached, it is placed on a random cell, until 20 markers are reached.
Task Reward Return is given by the number of reached markers divided by 20.
Appendix D Baselines and their hyperparameters
We compares to four baselines proposed by Trivedi et al. [57],  Liu et al. [37], and Carvalho et al. [10]. LEAPS, HPRL, and CEBS search programs in the latent spaces. HC search program in the programmatic space. The pseudo-code of CEBS is described in Algorithm 1, and the pseudo-code HC is described in Algorithm 2.
D.1 Learning Embeddings for Latent Program Synthesis (LEAPS)
The hyperparameters of the LEAPS baseline are from LEAPS [57] and HPRL [37]. We downloaded the pre-trained weights of LEAPS and used the parameters for searching programs. We do not use the official implementation of LEAPS, thus we do not have hyperparameters of exponential decay. Also, the unofficial implementation samples candidate from a fixed normal distribution which is hyperparameter in LEAPS [57]. In  Table LABEL:table:leaps_karel_setting, $K$ is the neighborhood size and $E$ is the candidate, $\sigma$ is the noise scale. In LEAPS and HPRL, they use an elite ratio to represent the candidate programs. For better understanding, we round $E$ to be an integer.
Table 7: LEAPS hyperparameter settings on all Karel tasks.
Task Name $K$ $\sigma$ $E$
StairClimber 32 0.25 2
FourCorners 64 0.5 13
TopOff 64 0.25 3
Maze 16 0.1 2
CleanHouse 32 0.25 2
Harvester 32 0.5 3
DoorKey 32 0.25 3
OneStroke 64 0.5 3
Seeder 32 0.25 3
Snake 32 0.25 6
D.2 Hierarchical programmatic reinforcement learning (HPRL)
HPRL [37] aims to improve LEAPS by composing several programs to represent more complex behaviors. Given a learned latent space, meta-policy learns to predict a sequence of actions, i.e. programs, by optimizing the return obtained by executing the composed programs using reinforcement learning. HPRL considers the discount factor $\gamma$ in the meta-MDP; as a result, the evaluation method is not the same as our problem formulation. To ensure a fair comparison, we record one million programs explored by meta-policy in the training stage. In HRPL, the authors use the training step as a hyperparameter. One training step is one step of meta-policy, and the maximum episode length is set to five as the original setting in the HPRL paper. In Table LABEL:table:hprl_karel_setting, we list the hyperparameters we modified in the HPRL training script. All other hyperparameters remain the same as described in HPRL.
Table 8: HRPL hyperparameter settings on all Karel tasks.
Task Name training steps height weight
StairClimber 50K 12 12
FourCorners 500K 12 12
TopOff 5M 12 12
Maze 50K 8 8
CleanHouse 500K 12 22
Harvester 5M 8 8
DoorKey 5M 8 8
OneStroke 5M 8 8
Seeder 5M 8 8
Snake 5M 8 8
D.3 Cross-entropy beam search (CEBS)
CEBS [10] also utilizes the pre-trained VAE weight from LEAPS [57] to search program in the latent space. All the hyperparameter follows the original CEBS with neighborhood $K$ equal to 64, candidate $E$ equal to 16, and noise ratio $\sigma$ equal to 0.25 for all ten Karel tasks. The pseudo-code is described in  Algorithm 1.
D.4 Hill climbing (HC)
HC [10] is a state-of-the-art algorithm solving Programmatic Reinforcement Learning tasks. All the hyperparameter follows the original HC with neighborhood $K$ equal to 250 for all ten Karel tasks. The pseudo-code is described in  Algorithm 2.
Algorithm 1 Cross-entropy beam search algorithm  [10]
0:  $K$, number of neighborhood; $E$ number of top candidate; $T$, the task; VAE, the program encoder and decoder.
0:  $\rho^{*}$ the highest averaged return over 32 task variants.
1:  $z\sim N(0,\mathbf{I})$
2:  $\rho\leftarrow\text{VAE.decode}(z)$
3:  $steps\leftarrow 0$
4:  $Return\leftarrow\text{evaluate}(\rho,T)$
5:  $\rho^{*}\leftarrow\rho$
6:  $Mean\leftarrow-\infty$
7:  $P\leftarrow\text{get-neighbor}(\rho,K)$
8:  while $steps<1000000$ do
9:     $Candidates\leftarrow\text{[]}$]
10:     for each $\rho_{\text{new}}$ in $P$ do
11:        $r\leftarrow\text{evaluate}(\rho_{\text{new}},T)$
12:        $Candidates.\text{append}(r,\rho_{\text{new}})$
13:        if $r>Return$ then
14:           $\rho^{*}\leftarrow\rho_{\text{new}}$
15:           $Return\leftarrow r$
16:        end if
17:     end for
18:     $Elites\leftarrow\text{Top-E}(Candidates)$
19:     if $Mean>\text{get-mean}(Elites)$ then
20:        break
21:     end if
22:     if $Return=1$ then
23:        break
24:     end if
25:     $Mean\leftarrow\text{get-mean}(Elites)$
26:     $P\leftarrow\text{[]}$]
27:     for  each $\rho_{\text{new}}$ in $Elites$ do
28:        $P.\text{extend}(\text{get-neighbor}(\rho_{\text{new}},K/E))$
29:     end for
30:     $steps\leftarrow steps+K$
31:  end while
Algorithm 2 Hill climbing algorithm  [10]
0:  $T$, the task; $K$, number of neighborhood.
0:  $\rho^{*}$, the highest averaged return over 32 task variants.
1:  Initialize $\rho$ with a random solution
2:  $Return\leftarrow\text{evaluate}(\rho,T)$
3:  $\rho^{*}\leftarrow\rho$
4:  $improved\leftarrow\text{True}$
5:  $steps\leftarrow 0$
6:  while $improved$ $steps<n$ do
7:     $improved\leftarrow\text{False}$
8:     $Neighbors\leftarrow\text{get-neighbor}(\rho,$K))))
9:     for each $\rho_{\text{new}}$ in $Neighbors$ do
10:        $r_{\text{new}}\leftarrow\text{evaluate}(\rho_{\text{new}},T)$
11:        if $r_{\text{new}}>Return$ then
12:           $\rho^{*}\leftarrow\rho_{\text{new}}$
13:           $Return\leftarrow r_{\text{new}}$
14:           $improved\leftarrow\text{True}$
15:           break
16:        end if
17:        $steps\leftarrow steps+1$
18:     end for
19:     $\rho\leftarrow\rho^{*}$
20:  end while
Appendix E An example pipeline of our method
This is our pipeline of getting programs from LLM with the task DoorKey.  Section E.1 details the system prompt,  Section E.2 provides all the information in one specific task,  Section E.3 lists the LLM response, and  Section E.4 gives an example how we get the Karel DSL program.
E.1 The system prompt
Youre currently navigating within a Karel environment, which is essentially a grid world. In this context, a "world" is referred to as a "map." Within this map, theres an entity known as the "agent," capable of movement, changing direction, as well as picking up and placing markers on the map. Additionally, there are obstacles called "walls" that impede the agents progress; whenever the agent encounters a wall, it turns around. Furthermore, there are pre-existing "markers" scattered throughout the map at the beginning, though the agent has the ability to both pickup and place these markers as needed.
Your objective is to generate the appropriate Python program based on a given task name and description. This Python program will encompass actions enabling the agent to engage with the environment, alongside perceptions facilitating the agents recognition of the environments dynamics.
Here are the available actions for the agent:
move(): Asks the agent to move forward one cell. The agent will instead turn left twice if a wall is blocking its way.
turnLeft(): Asks the agent to rotate 90 degrees counterclockwise.
turnRight(): Asks the agent to rotate 90 degrees clockwise.
pickMarker(): Asks the agent to pick up one marker from the current cell.
putMarker(): Asks the agent to put down one marker on the current cell.
Here are the available perceptions of the agent:
frontIsClear(): Returns True if there is no wall in front of the agent.
leftIsClear(): Returns True if there is no wall on the agents left.rightIsClear(): Returns True if there is no wall on the agents right.
markersPresent(): Returns True if there exist markers on the current cell.
noMarkersPresent(): Returns True if there is no marker on the current cell.
There are some limitations for the Python program:
- do not define other functions besides run()
- do not call other functions
- do not define variables
- do not use True, False, break, continue, return, ==, !=, elif, or, and
Python to Karel dsl conversion
1. "def run(): s" to "DEF run m( s m)"
2. "while b: s" to "WHILE c( b c) w( s w)"
3. "if b: s" to "IF c( b c) i( s I)"
4. "if b: s else: s" to "IFELSE c( b c) i( s i) ELSE e( s e)"
5. "for i in range(n): s" to "REPEAT R=n r( s r)"
6. "not h" to "not c( h c)"
7. "frontIsClear()" to "frontIsClear"
8. "leftIsClear()" to "leftIsClear"
9. "rightIsClear()" to "rightIsClear"
10. "markersPresent()" to "markersPresent"
11. "noMarkersPresent()" to "noMarkersPresent"
12. "move()" to "move"
13. "turnLeft()" to "turnLeft"
14. "turnRight()" to "turnRight"
15. "putMarker()" to "putMarker"
16. "pickMarker()" to "pickMarker"
E.2 The user prompt of the task DoorKey
Ill provide you with the task name and description.
Task name: DOORKEY
Task map: The map is a 8x8 grid surrounded by walls that is vertically split into two chambers. The left chamber is 6x3 grid and the right chamber is 6x2 grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Task agent position: The agent starts on a random cell on the left chamber facing east.
Task goal: The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
Task return: Picking up the first marker yields a 0.5 reward, and putting a marker on the goal marker yields an additional 0.5.
1. Generate 1 simple and short Python program to tackle the task, avoid using comments.
2. Convert the Python program to the Karel dsl program.
E.3 The sample response of LLM in the task DoorKey
### Python Program for Task DOORKEY
‘‘‘python
def run():
while not markersPresent():
if frontIsClear():
move()
else:
turnLeft()
pickMarker()
turnLeft()
while not markersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
### Karel DSL Conversion
DEF run m(
WHILE not c( markersPresent c) w(
IF frontIsClear i(
move i)
ELSE e(
turnLeft e) w)
pickMarker m
turnLeft m
WHILE not c( markersPresent c) w(
IF frontIsClear i(
move i)
ELSE e(
turnLeft e) w)
putMarker m)
E.4 Apply post-processing and retrieve the final program
For the Python program part in LLM response, we convert the Python program to a DSL program through the rule-based method and apply post-processing to correct minor mistakes. For the DSL part, we directly apply post-processing to correct minor mistakes. Section E.4.1 presents the original Python program, the converted DSL, and the post-processed DSL. Section E.4.2 shows the original DSL and the post-processed DSL. All of the post-processing rules are listed in Table 5.
E.4.1 Python program part
Listing 1: Original Python program
def run():
while not markersPresent():
if frontIsClear():
move()
else:
turnLeft()
pickMarker()
turnLeft()
while not markersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
Listing 2: DSL program converted through the rule-based method
DEF run m(
WHILE c( not c( markersPresent c) c) w(
IFELSE c( frontIsClear c) i(
move i)
ELSE e(
turnLeft e) w)
pickMarker
turnLeft
WHILE c( not c( markersPresent c) c) w(
IFELSE c( frontIsClear c) i(
move i)
ELSE e(
turnLeft e) w)
putMarker m)
Listing 3: DSL program after post-processing
DEF run m(
WHILE c( noMarkersPresent c) w(
IFELSE c( frontIsClear c) i(
move i)
ELSE e(
turnLeft e) w)
pickMarker
turnLeft
WHILE c( noMarkersPresent c) w(
IFELSE c( frontIsClear c) i(
move i)
ELSE e(
turnLeft e) w)
putMarker m)
E.4.2 DSL program part
Listing 4: Original DSL program
DEF run m(
WHILE not c( markersPresent c) w(
IF frontIsClear i(
move i)
ELSE e(
turnLeft e) w)
pickMarker m
turnLeft m
WHILE not c( markersPresent c) w(
IF frontIsClear i(
move i)
ELSE e(
turnLeft e) w)
putMarker m)
Listing 5: DSL program after post-processing
DEF run m(
WHILE c( noMarkersPresent c) w(
IFELSE c( frontIsClear c) i(
move i)
ELSE e(
turnLeft e) w)
pickMarker
turnLeft
WHILE c( noMarkersPresent c) w(
IFELSE c( frontIsClear c) i(
move i)
ELSE e(
turnLeft e) w)
putMarker m)
Appendix F LLM prompts for ablations and revision
Ablation prompts for program generation methods. We conduct an ablation study in Section 5.3 to justify that our LLM-generating program method performs best in acceptance rate and best return. Here are the complete prompts. There are three approaches to generating DSL programs: Pythonic-DSL, Python, and DSL. There are two types of prompt, system prompt and user prompt, in all approaches. In the system prompt, both the Python and the Pythonic-DSL approaches contain limitations of Python usage. The DSL approach contains the grammar of the Karel, and the Pythonic-DSL contains the paired Python-like and Karel production rules. All of the system prompts contain the environment physics, action, and perception. On the other hand, the user prompt contains five placeholders for task name, map position, initial position, task goal, and task reward. In Appendix C, all of the task-dependent information can be filled in the placeholders. Section E.2 shows the full Pythonic-DSL prompt with the task DoorKey filled in the placeholders. The Pythonic-DSL prompts are in  Section F.1, and the Python prompts are in Section F.2, and the DSL prompts are in Section F.3.
Prompts for LLM revision. We list all of the user prompts in the experiment of LLM revision, the system prompt is the same as the one in Section F.1. There are four approaches to revising the program: Regenerate, Regenerate with reward, Agent execution trace, and Agent and program execution trace. We ask the LLM to regenerate with all generated programs in the last round without repetition, and the Regenerate prompt is in Section F.4. We ask the LLM to regenerate with the program relating to reward, and the revision prompt of the Regenerate with reward is in Section F.5. We utilize the execution trace of the Karel agent in the grid world, and the revision prompt of the Agent execution trace is in Section F.6. We provided the action/perception call and executing line, and the revision prompt of the Agent and program execution trace is in Section F.7. The program Figure 11 is the revision target of our method revision method in Section F.6 and Section F.7. Figure 11 is the trajectories of the program we used in method Section F.6 and Section F.7. In the example, the origin program has an average return of 0.5. LLM revision with agent trace can reach 0.640625, and LLM revision with both agent and program traces can reach a great result of 0.8125.
(a) This is the initial state of one of DoorKey the task variants.
(b) This is the last step of the trajectories in the prompts.
(c) The program ends at step 47 and no reward is granted.
Figure 11: This is the demonstration of the task DoorKey. The Karel agent can only reach the surrounding grids at the left chamber in the trajectory.
{listing}
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() if frontIsClear(): move() turnRight() pickMarker() while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() putMarker() This is the program for LLM revision.
{listing}
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while frontIsClear(): move() turnRight() while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() if frontIsClear(): move() turnRight() putMarker() The revision results of Section F.6.
{listing}
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while noMarkersPresent(): if rightIsClear(): turnRight() move() else: if frontIsClear(): move() else: turnLeft() putMarker() The revision results of Section F.7.
F.1 Pythonic-DSL
F.1.1 System prompt
Youre currently navigating within a Karel environment, which is essentially a grid world. In this context, a "world" is referred to as a "map." Within this map, theres an entity known as the "agent," capable of movement, changing direction, as well as picking up and placing markers on the map. Additionally, there are obstacles called "walls" that impede the agents progress; whenever the agent encounters a wall, it turns around. Furthermore, there are pre-existing "markers" scattered throughout the map at the beginning, though the agent has the ability to both pickup and place these markers as needed.
Your objective is to generate the appropriate Python program based on a given task name and description. This Python program will encompass actions enabling the agent to engage with the environment, alongside perceptions facilitating the agents recognition of the environments dynamics.
Here are the available actions for the agent:
move(): Asks the agent to move forward one cell. The agent will instead turn left twice if a wall is blocking its way.
turnLeft(): Asks the agent to rotate 90 degrees counterclockwise.
turnRight(): Asks the agent to rotate 90 degrees clockwise.
pickMarker(): Asks the agent to pick up one marker from the current cell.
putMarker(): Asks the agent to put down one marker on the current cell.>
Here are the available perceptions of the agent:
frontIsClear(): Returns True if there is no wall in front of the agent.
leftIsClear(): Returns True if there is no wall on the agents left.
rightIsClear(): Returns True if there is no wall on the agents right.
markersPresent(): Returns True if there exist markers on the current cell.
noMarkersPresent(): Returns True if there is no marker on the current cell.
There are some limitations for the Python program:
- do not define other functions besides run()
- do not call other functions
- do not define variables
- do not use True, False, break, continue, return, ==, !=, elif, or, and
Python to Karel dsl conversion
1. "def run(): s" to "DEF run m( s m)"
2. "while b: s" to "WHILE c( b c) w( s w)"
3. "if b: s" to "IF c( b c) i( s i)"
4. "if b: s else: s" to "IFELSE c( b c) i( s i) ELSE e( s e)"
5. "for i in range(n): s" to "REPEAT R=n r( s r)"
6. "not h" to "not c( h c)"
7. "frontIsClear()" to "frontIsClear"
8. "leftIsClear()" to "leftIsClear"
9. "rightIsClear()" to "rightIsClear"
10. "markersPresent()" to "markersPresent"
11. "noMarkersPresent()" to "noMarkersPresent"
12. "move()" to "move"
13. "turnLeft()" to "turnLeft"
14. "turnRight()" to "turnRight"
15. "putMarker()" to "putMarker"
16. "pickMarker()" to "pickMarker"
F.1.2 User prompt
Ill provide you with the task name and description.
Task name: <<Task Name>>
Task map: <<Map Description>>
Task agent position: <<Initial Position>>
Task goal: <<Task Goal>>
Task return: <<Task Reward>>
1. Generate 1 simple and short Python program to tackle the task, avoid using comments.
2. Convert the Python program to the Karel dsl program.
F.2 Python
F.2.1 System prompt
Youre currently navigating within a Karel environment, which is essentially a grid world. In this context, a "world" is referred to as a "map." Within this map, theres an entity known as the "agent," capable of movement, changing direction, as well as picking up and placing markers on the map. Additionally, there are obstacles called "walls" that impede the agents progress; whenever the agent encounters a wall, it turns around. Furthermore, there are pre-existing "markers" scattered throughout the map at the beginning, though the agent has the ability to both pickup and place these markers as needed.
Your objective is to generate the appropriate Python program based on a given task name and description. This Python program will encompass actions enabling the agent to engage with the environment, alongside perceptions facilitating the agents recognition of the environments dynamics.
Here are the available actions for the agent:
move(): Asks the agent to move forward one cell. The agent will instead turn left twice if a wall is blocking its way.
turnLeft(): Asks the agent to rotate 90 degrees counterclockwise.
turnRight(): Asks the agent to rotate 90 degrees clockwise.
pickMarker(): Asks the agent to pick up one marker from the current cell.
putMarker(): Asks the agent to put down one marker on the current cell.
Here are the available perceptions of the agent:
frontIsClear(): Returns True if there is no wall in front of the agent.
leftIsClear(): Returns True if there is no wall on the agents left.
rightIsClear(): Returns True if there is no wall on the agents right.
markersPresent(): Returns True if there exist markers on the current cell.
noMarkersPresent(): Returns True if there is no marker on the current cell.
There are some limitations for the Python program:
- do not define other functions besides run()
- do not call other functions
- do not define variables
- do not use True, False, break, continue, return, ==, !=, elif, or, and
F.2.2 User prompt
Ill provide you with the task name and description.
Task name: <<Task Name>>
Task map: <<Map Description>>
Task agent position: <<Initial Position>>
Task goal: <<Task Goal>>
Task return: <<Task Reward>>
1. Generate 1 simple and short Python program to tackle the task, avoid using comments.
F.3 DSL
F.3.1 System prompt
Youre currently navigating within a Karel environment, which is essentially a grid world. In this context, a "world" is referred to as a "map." Within this map, theres an entity known as the "agent," capable of movement, changing direction, as well as picking up and placing markers on the map. Additionally, there are obstacles called "walls" that impede the agents progress; whenever the agent encounters a wall, it turns around. Furthermore, there are pre-existing "markers" scattered throughout the map at the beginning, though the agent has the ability to both pickup and place these markers as needed.
Your objective is to generate the appropriate Karel dsl program based on a given task name and description. This Karel dsl program will encompass actions enabling the agent to engage with the environment, alongside perceptions facilitating the agents recognition of the environments dynamics.
Here are the available actions for the agent:
move: Asks the agent to move forward one cell. The agent will instead turn left twice if a wall is blocking its way.
turnLeft: Asks the agent to rotate 90 degrees counterclockwise.
turnRight: Asks the agent to rotate 90 degrees clockwise.
pickMarker: Asks the agent to pick up one marker from the current cell.
putMarker: Asks the agent to put down one marker on the current cell.
Here are the available perceptions of the agent:
frontIsClear: Returns True if there is no wall in front of the agent.
leftIsClear: Returns True if there is no wall on the agents left.
rightIsClear: Returns True if there is no wall on the agents right.
markersPresent: Returns True if there exist markers on the current cell.
noMarkersPresent: Returns True if there is no marker on the current cell.
This is the production role of the domain-specific language of the Karel environment.
Program p := DEF run m( s m)
Statement s := WHILE c( b c) w( s w) | IF c( b c) i( s i) | IFELSE c( b c) i( s i) ELSE e( s e) | REPEAT R=n r( s r) | s s | a
Condition b := h | not c( h c)
Number n := 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19
Perception h := frontIsClear | leftIsClear | rightIsClear | markersPresent | noMarkersPresent
Action a := move | turnLeft | turnRight | putMarker | pickMarker
F.3.2 User prompt
Ill provide you with the task name and description.
Task name: <<Task Name>>
Task map: <<Map Description>>
Task agent position: <<Initial Position>>
Task goal: <<Task Goal>>
Task return: <<Task Reward>>
1. Generate 1 simple and short Karel dsl program to tackle the task, avoid using comments.
F.4 Regenerate
Ill provide you with the task name, task description, and the programs you generated last time.
Task name: DOORKEY
Task map: The map is a 8x8 grid surrounded by walls that is vertically split into two chambers. The left chamber is 6x3 grid and the right chamber is 6x2 grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Task agent position: The agent starts on a random cell on the left chamber facing east.
Task goal: The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
Task return: Picking up the first marker yields a 0.5 reward, and putting a marker on the goal marker yields an additional 0.5.
These are the programs you generated last time, all of these programs cannot yield perfect performance.
Program 1:
def run():
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
if markersPresent():
putMarker()
###23 programs are truncated.###
Program 25:
def run():
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
pickMarker()
for i in range(19):
if frontIsClear():
move()
else:
turnRight()
1. Generate a Python program that is not identical to any of the previous programs to tackle the task, and avoid using comments.
2. Convert the Python program to the Karel dsl program.
F.5 Regenerate with reward
Ill provide you with the task name, task description, and the programs rewards pairs sorted by their evaluation rewards from 32 task variants.
Task name: DOORKEY
Task map: The map is a 8x8 grid surrounded by walls that is vertically split into two chambers. The left chamber is 6x3 grid and the right chamber is 6x2 grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Task agent position: The agent starts on a random cell on the left chamber facing east.
Task goal: The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
Task return: Picking up the first marker yields a 0.5 reward, and putting a marker on the goal marker yields an additional 0.5.
Program reward pairs sorted by their evaluation rewards:
Program 1:
def run():
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
reward:
0.5
###23 programs are truncated.###
Program 25:
def run():
for i in range(19):
if markersPresent():
pickMarker()
while not frontIsClear():
turnLeft()
move()
if frontIsClear():
move()
if markersPresent():
putMarker()
reward:
-0.5
1. Depending on this information, examine the program pattern that the highest score programs process, but the lowest score programs do not.
2. Generate 1 simple and short Python program according to the pattern to tackle the task, avoid using comment.
3. Convert the Python program to the Karel dsl program.
F.6 Agent execution trace
Ill provide you with the code you developed previously, with the goal of refining it. To guide your revision, youll receive the specific task name and a description. Since there are 32 different versions of the task that share the same objective but differ by random seeds, I will identify the specific variant where the performance of the program is most lacking. Additionally, youll get the initial state of the task, the code, and a detailed trajectory demonstrating how the code operates within this particular scenario. This trajectory will detail each action step-by-step and show a localized snapshot of the environment (a 3x3 area centered on the agent) during execution. Rewards received by the agent will also be shown during these steps.
Task name: DOORKEY
Task map: The map is a 8x8 grid surrounded by walls that is vertically split into two chambers. The left chamber is 6x3 grid and the right chamber is 6x2 grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Task agent position: The agent starts on a random cell on the left chamber facing east.
Task goal: The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
Task return: Picking up the first marker yields a 0.5 reward, and putting a marker on the goal marker yields an additional 0.5.
Initial state:
Wall(0, 0) ; Wall(0, 1) ; Wall(0, 2) ; Wall(0, 3) ; Wall(0, 4) ; Wall(0, 5) ; Wall(0, 6) ; Wall(0, 7) ;
Wall(1, 0) ; Empty(1, 1) ; Empty(1, 2) ; Empty(1, 3) ; Wall(1, 4) ; Empty(1, 5) ; Marker(1, 6, quantity=1) ; Wall(1, 7) ;
Wall(2, 0) ; Empty(2, 1) ; Empty(2, 2) ; Empty(2, 3) ; Wall(2, 4) ; Empty(2, 5) ; Empty(2, 6) ; Wall(2, 7) ;
Wall(3, 0) ; Empty(3, 1) ; Empty(3, 2) ; Empty(3, 3) ; Wall(3, 4) ; Empty(3, 5) ; Empty(3, 6) ; Wall(3, 7) ;
Wall(4, 0) ; Empty(4, 1) ; Agent(4, 2, direction=(0, 1)) ; Empty(4, 3) ; Wall(4, 4) ; Empty(4, 5) ; Empty(4, 6) ; Wall(4, 7) ;
Wall(5, 0) ; Empty(5, 1) ; Marker(5, 2, quantity=1) ; Empty(5, 3) ; Wall(5, 4) ; Empty(5, 5) ; Empty(5, 6) ; Wall(5, 7) ;
Wall(6, 0) ; Empty(6, 1) ; Empty(6, 2) ; Empty(6, 3) ; Wall(6, 4) ; Empty(6, 5) ; Empty(6, 6) ; Wall(6, 7) ;
Wall(7, 0) ; Wall(7, 1) ; Wall(7, 2) ; Wall(7, 3) ; Wall(7, 4) ; Wall(7, 5) ; Wall(7, 6) ; Wall(7, 7) ;
Program:
def run():
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
The average reward on 32 task variants is:
0.5
Trajectory:
Step 1:
Agent performs a perception: noMarkersPresent. The result is True.
Partial state:
Empty(3, 1) ; Empty(3, 2) ; Empty(3, 3) ;
Empty(4, 1) ; Agent(4, 2, direction=(0, 1)) ; Empty(4, 3) ;
Empty(5, 1) ; Marker(5, 2, quantity=1) ; Empty(5, 3) ;
Step 2:
Agent performs a perception: frontIsClear. The result is True.
Partial state:
Empty(3, 1) ; Empty(3, 2) ; Empty(3, 3) ;
Empty(4, 1) ; Agent(4, 2, direction=(0, 1)) ; Empty(4, 3) ;
Empty(5, 1) ; Marker(5, 2, quantity=1) ; Empty(5, 3) ;
Step 3:
Agent performs an action: move.
Partial state:
Empty(3, 2) ; Empty(3, 3) ; Wall(3, 4) ;
Empty(4, 2) ; Agent(4, 3, direction=(0, 1)) ; Wall(4, 4) ;
Marker(5, 2, quantity=1) ; Empty(5, 3) ; Wall(5, 4) ;
###45 steps are truncated.###
Step 49:
Agent performs a perception: frontIsClear. The result is False.
Partial state:
Wall(1, 0) ; Empty(1, 1) ; Empty(1, 2) ;
Wall(2, 0) ; Agent(2, 1, direction=(0, -1)) ; Empty(2, 2) ;
Wall(3, 0) ; Empty(3, 1) ; Empty(3, 2) ;
The total step number is 105, the latter ones are truncated.
The total reward is 0.0
1. Depending on this information, please analyze the reason why the program failed to achieve 1.0 on this task variant and generate a new strategy to solve this task.
2. Generate 1 simple and short Python program according to the new strategy to tackle the task, avoid using comment.
3. Convert the Python program to the Karel dsl program.
F.7 Agent and Program execution trace.
Ill provide you with the code you developed previously, with the goal of refining it. To guide your revision, youll receive the specific task name and a description. Since there are 32 different versions of the task that share the same objective but differ by random seeds, I will identify the specific variant where the performance of the program is most lacking. Additionally, youll get the initial state of the task, the code, and a detailed trajectory demonstrating how the code operates within this particular scenario. This trajectory will detail each action step-by-step, indicate which section of your code is active, and show a localized snapshot of the environment (a 3x3 area centered on the agent) during execution. Rewards received by the agent will also be shown during these steps.
Task name: DOORKEY
Task map: The map is a 8x8 grid surrounded by walls that is vertically split into two chambers. The left chamber is 6x3 grid and the right chamber is 6x2 grid. There is a marker placed randomly on the left chamber as a key, and another marker placed randomly on the right chamber as a goal.
Task agent position: The agent starts on a random cell on the left chamber facing east.
Task goal: The goal of the agent is to pick up a marker on the left chamber, which opens a door connecting both chambers. Allow the agent to reach and put a marker on the goal marker.
Task return: Picking up the first marker yields a 0.5 reward, and putting a marker on the goal marker yields an additional 0.5.
Initial state:
Wall(0, 0) ; Wall(0, 1) ; Wall(0, 2) ; Wall(0, 3) ; Wall(0, 4) ; Wall(0, 5) ; Wall(0, 6) ; Wall(0, 7) ;
Wall(1, 0) ; Empty(1, 1) ; Empty(1, 2) ; Empty(1, 3) ; Wall(1, 4) ; Empty(1, 5) ; Marker(1, 6, quantity=1) ; Wall(1, 7) ;
Wall(2, 0) ; Empty(2, 1) ; Empty(2, 2) ; Empty(2, 3) ; Wall(2, 4) ; Empty(2, 5) ; Empty(2, 6) ; Wall(2, 7) ;
Wall(3, 0) ; Empty(3, 1) ; Empty(3, 2) ; Empty(3, 3) ; Wall(3, 4) ; Empty(3, 5) ; Empty(3, 6) ; Wall(3, 7) ;
Wall(4, 0) ; Empty(4, 1) ; Agent(4, 2, direction=(0, 1)) ; Empty(4, 3) ; Wall(4, 4) ; Empty(4, 5) ; Empty(4, 6) ; Wall(4, 7) ;
Wall(5, 0) ; Empty(5, 1) ; Marker(5, 2, quantity=1) ; Empty(5, 3) ; Wall(5, 4) ; Empty(5, 5) ; Empty(5, 6) ; Wall(5, 7) ;
Wall(6, 0) ; Empty(6, 1) ; Empty(6, 2) ; Empty(6, 3) ; Wall(6, 4) ; Empty(6, 5) ; Empty(6, 6) ; Wall(6, 7) ;
Wall(7, 0) ; Wall(7, 1) ; Wall(7, 2) ; Wall(7, 3) ; Wall(7, 4) ; Wall(7, 5) ; Wall(7, 6) ; Wall(7, 7) ;
Program:
def run():
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
The average reward on 32 task variants is:
0.5
Trajectory:
Step 1:
Program:
def run():
while noMarkersPresent(): # Currently executing this line
if frontIsClear():
move()
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
Agent performs a perception: noMarkersPresent. The result is True.
Partial state:
Empty(3, 1) ; Empty(3, 2) ; Empty(3, 3) ;
Empty(4, 1) ; Agent(4, 2, direction=(0, 1)) ; Empty(4, 3) ;
Empty(5, 1) ; Marker(5, 2, quantity=1) ; Empty(5, 3) ;
Step 2:
Program:
def run():
while noMarkersPresent():
if frontIsClear(): # Currently executing this line
move()
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
Agent performs a perception: frontIsClear. The result is True.
Partial state:
Empty(3, 1) ; Empty(3, 2) ; Empty(3, 3) ;
Empty(4, 1) ; Agent(4, 2, direction=(0, 1)) ; Empty(4, 3) ;
Empty(5, 1) ; Marker(5, 2, quantity=1) ; Empty(5, 3) ;
Step 3:
Program:
def run():
while noMarkersPresent():
if frontIsClear():
move() # Currently executing this line
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
Agent performs an action: move.
Partial state:
Empty(3, 2) ; Empty(3, 3) ; Wall(3, 4) ;
Empty(4, 2) ; Agent(4, 3, direction=(0, 1)) ; Wall(4, 4) ;
Marker(5, 2, quantity=1) ; Empty(5, 3) ; Wall(5, 4) ;
###45 steps are truncated.###
Step 49:
Program:
def run():
while noMarkersPresent():
if frontIsClear(): # Currently executing this line
move()
else:
turnLeft()
if frontIsClear():
move()
turnRight()
pickMarker()
while noMarkersPresent():
if frontIsClear():
move()
else:
turnLeft()
putMarker()
Agent performs a perception: frontIsClear. The result is False.
Partial state:
Wall(1, 0) ; Empty(1, 1) ; Empty(1, 2) ;
Wall(2, 0) ; Agent(2, 1, direction=(0, -1)) ; Empty(2, 2) ;
Wall(3, 0) ; Empty(3, 1) ; Empty(3, 2) ;
The total step number is 105, the latter ones are truncated.
The total reward is 0.0
1. Depending on this information, please analyze the reason why the program failed to achieve 1.0 on this task variant and generate a new strategy to solve this task.
2. Generate 1 simple and short Python program according to the new strategy to tackle the task, avoid using comment.
3. Convert the Python program to the Karel dsl program.